√64-√p+√q=9 find p+q
1
Expert's answer
2013-03-06T05:28:00-0500
√64-√p+√q=9 find p+q
-√p+√q=9-8
-√p+√q=1
Moreover p and q are natural numbers.
After Squaring we have: p+q-2*sqrt(pq)=1
Then q=9, p=4.